# LVLM-Count: Enhancing the Counting Ability of Large Vision-Language Models

### v3

@@ 4.2 Results
Emoji-Count. We evaluate the performance of the LVLMs and their performance using LVLM-Count on the Emoji-Count benchmark. The results are shown in Table 3. This is a challenging benchmark, as it requires understanding complex concepts. We observe that taking the number of masks as the final count performs particularly poorly, as for any object of interest in the image, the segmentation stage tends to segment all the objects and cannot distinguish between different icons. Although GPT-4o and Gemma 3 show reasonable performance, the other open-source model does not perform well. Nonetheless, the performance of all three base LVLMs is significantly enhanced by LVLM-Count, especially Qwen2, which performs almost on par with GPT-4o when LVLM-Count is used for it.
Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ Number of the target segmentation masks 82.47 - 107.98 - GPT-4o counting target segmentaion masks 107.72 - 162.12 - GPT-4o 22.51 - 35.94 - LVLM-Count (GPT-4o as LVLM) 16.29 $\downarrow 6.22$ 32.47 $\downarrow 3.47$ Gemma 3 27B 21.43 - 24.15 - LVLM-Count (Gemma3 27B as LVLM) 15.76 $\downarrow 5.67$ 21.00 $\downarrow 3.15$ Qwen2 VL 72B AWQ 82.41 - 186.32 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 23.74 $\downarrow 58.67$ 40.04 $\downarrow 143.84$
@@
Penguin Benchmark. We report the counting performance on the Penguin benchmark in Table 4. We evaluate two variants of LVLM-Count. The primary variant uses GroundingDINO in both the area detection and target segmentation stages. The alternative variant does not use a detection model; instead, SAM is configured to segment all entities in the image. Both variants improve MAE across all three LVLMs, highlighting LVLM-Count’s robustness in scenarios with heavy occlusion and complex backgrounds—conditions that pose significant challenges for area detection and target segmentation. Additional details and a visual example showing the segmentation masks and division paths generated in both variants for a sample from the Penguin benchmark can be found in Appendix D.
Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ GPT4o 36.01 - 46.64 - LVLM-Count( Main variant, using GPT-4o ) 26.95 $\downarrow$ 9.06 39.32 $\downarrow$ 7.32 LVLM-Count(SAM-only variant, using GPT-4o) 28.79 $\downarrow$ 7.22 45.71 $\downarrow$ 0.93 Gemma 3 27B 49.87 - 61.06 - LVLM-Count( Main variant, using Gemma 3 ) 36.75 $\downarrow$ 13.12 47.24 $\downarrow$ 13.82 LVLM-Count( SAM-only variant, using Gemma 3 ) 41.97 $\downarrow$ 7.90 54.55 $\downarrow$ 6.51 Qwen2 VL 72B AWQ 44.96 - 66.78 - LVLM-Count( Main variant, using Qwen2 ) 28.67 $\downarrow$ 16.29 43.71 $\downarrow$ 23.07 LVLM-Count(SAM-only, using Qwen2 ) 38.29 $\downarrow$ 6.67 51.47 $\downarrow$ 15.31
@@
Limitations
@@ Appendix A Ablation study
We perform ablation studies using GPT-4o, conducting experiments on the FSC-147 test set. Table 5 presents the results of these ablation experiments.
Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ GPT-4o 25.17 - 137.86 - GPT-4o + Naive division 33.04 $\uparrow$ 7.87 116.83 $\downarrow$ 21.03 GPT-4o + Area Detection + Naive division 32.69 $\uparrow$ 7.52 102.41 $\downarrow$ 35.45 GPT-4o + Area detection 23.08 $\downarrow$ 2.09 120.06 $\downarrow$ 17.8 GPT-4o + Object-aware division (using SAM without GroundingDINO) 21.01 $\downarrow$ 4.16 135.04 $\downarrow$ 2.82 GPT-4o + Object-aware division 19.17 $\downarrow$ 6.00 120.61 $\downarrow$ 17.25 GPT-4o + Area detection + Object-aware division, (equiv. to LVLM-Count) 18.09 $\downarrow$ 7.08 88.77 $\downarrow$ 49.09
@@
Figure 8 illustrates three examples of LVLM-Count’s performance on FSC-147. Additionally, Figure 9 includes an example from the PASCAL VOC benchmark. Finally, we present three visual examples from the Emoji-Count benchmark in Figure 10. In these visual examples, LVLM-Count achieves more accurate results compared to the base GPT-4o.
@@ Appendix C Robustness to Crowded Scenes and Heavy Occlusion
Figure 11 demonstrates the strong performance of our method on a randomly selected image from the web containing various types of flowers. This scene features extreme occlusion; nonetheless, our method effectively handles the division task. It is important to note that the strong performance of our method in this challenging scenario is not coincidental. Figure 12 illustrates the impact of our post-processing operations. The upper row displays the unprocessed SAM masks, while the bottom row shows the results after post-processing, which generates reliable paths for the division lines.
@@
To further demonstrate the effectiveness of our method in crowded scenes, we present Table 6, which compares the performance of different LVLMs on FSC-147 dataset samples where the ground truth count is 100 or higher. The table contrasts their baseline performance with their performance when using our LVLM-Count method. The results show significant improvements when LVLM-Count is employed. Furthermore, Figure 13 presents three visual examples of LVLM-Count’s performance on crowded images from FSC-147.
@@ Appendix D Additional Information and Visual Examples from the Penguin Benchmark
We construct our benchmark using the separate-site split. Given the dataset’s substantial size, which comprises tens of thousands of images, we randomly select 100 samples from the chosen split. To ensure balance, the sampling probability for an image with a specific ground-truth annotation count is inversely proportional to the frequency of that count in the entire split. We evaluate two variants of our pipeline. The main variant employs GroundingDINO with a low threshold for area detection and target segmentation, while the alternative variant does not use GroundingDINO at all. In this variant, masks are obtained by configuring SAM to segment any entity in the image. The results presented in Table 4 show that both variants significantly improve the counting performance of the LVLMs, demonstrating the robustness of the initial stages in our pipeline.
Figure 14 demonstrates the segmentation masks and division paths for the two variants on a sample from the Penguin benchmark. In Figure 14 (a), GroundingDINO with a low detection threshold is employed. It correctly identifies all penguin instances, though there are some false positives. These detections are subsequently segmented by SAM. Figure 14 (c) depicts the division paths found using these masks. In Figure 14 (b), GroundingDINO is not used, and SAM is configured to segment all entities in the scene. As a result, in addition to the penguins, a large number of other objects are segmented. Nevertheless, Figure 14 (d) shows that the division paths derived from these masks successfully partition the image without cutting any penguins. Note that some excessively large masks generated by SAM have been removed due to post-processing in our pipeline.
@@ Appendix E Real-world Application of LVLM-Count
As stated in Section 1, counting has numerous real-world applications, including but not limited to biology, health, industry, warehousing, and environmental monitoring. Below, we demonstrate the performance of LVLM-Count on examples from the following areas: i) biology/health, ii) industry/warehousing, and iii) environmental monitoring. We also compare its results with those of the base LVLM (GPT-4o for the figures in this section). Note that in all examples, the cluster-based approach automatically determines the start and end points of the division paths.
In Figure 15, images of two laboratory samples are analyzed using LVLM-Count. The first row shows an image from a dataset introduced by Lempitsky and Zisserman [2010], which contains simulated bacterial cells from fluorescence-light microscopy, created by Lehmussola et al. [2007]. The second row shows an image from the BM dataset introduced by Kainz et al. [2015], which contains bone marrow samples from eight healthy individuals. The standard staining procedure highlights the nuclei of various cell types in blue, while other cellular components appear in shades of pink and red [Paul Cohen et al., 2017]. As observed, LVLM-Count achieves much higher accuracy in counting bacterial cells and bone marrow nuclei in the top and bottom rows of Figure 15, respectively, compared to the base LVLM, particularly for the bone marrow nuclei.
In Figure 16, two images from industrial scenes are analyzed using LVLM-Count. The top row shows a sectional image of a stockpile of tree logs, and the bottom row shows an image from an industrial area containing barrels of various colors. For the top image, the objects of interest are the tree logs, while for the bottom image, LVLM-Count is tasked with counting the blue barrels. In both cases, LVLM-Count’s predictions are significantly closer to the ground truth values than those of the base LVLM, particularly for the tree logs, where the ground truth number is too large for the base LVLM to estimate accurately.
Figure 17 shows an image sourced from a dataset [Penguin Research, 2016] created as part of an ongoing initiative to monitor the penguin population in Antarctica. This dataset comprises images captured hourly by a network of fixed cameras installed at over 40 locations. Over several years, this effort has accumulated over 500,000 images. Zoologists use these images to identify trends in penguin population sizes at each site, facilitating studies on potential correlations with factors such as climate change. Thus, determining the number of penguins in each image is crucial. Given the challenges of engaging human annotators to process such a vast dataset, automating the counting task is highly desirable [Arteta et al., 2016]. LVLM-Count is prompted to count the number of penguins in the image, and as observed, its predictions are significantly closer to the ground truth than those of the base LVLM.
@@ Appendix F LVLM-Count’s Power in Handling Multiple Object Categories in the Same Image
The image in Figure 18 contains three object categories: person, cow, and horse. In the top row, the object of interest is “cow.” We prompt LVLM-Count to count the cows. First, the masks are produced through the initial stages of our pipeline, and then the cluster-based approach is used to automatically determine the start and end points of the division paths. It can be observed that horses have also been masked as cows. Nonetheless, this does not negatively impact the final answer; it merely causes the division lines to avoid cutting through the horses as well. The counting in LVLM-Count is performed by an LVLM (GPT-4o in this figure) and does not rely on the masks. We observe that GPT-4o successfully counts the number of cows in the resulting subimages, leading to the correct final answer.
In the middle row of Figure 18, the object of interest is “person.” LVLM-Count again successfully counts the number of people accurately. A more interesting case is the bottom row of Figure 18, where both cows and persons are objects of interest. We prompt LVLM-Count to count the number of “cows and persons.” Similar to the first row of the figure, there are false positive masks here as well. However, LVLM-Count successfully counts the number of instances from both categories combined since the counting is ultimately performed by the LVLM. Note that the number of objects in this image is limited, and GPT-4o might answer these questions correctly without the need for the LVLM-Count pipeline. This image has been chosen to illustrate LVLM-Count’s power in handling multiple objects in a counting task rather than for comparison with the baseline LVLM.
@@ Appendix G False Positive Masks at the Target Segmentation Stage
For example, in the top row of Figure 19, $E=$brown egg”. However, all the eggs have been segmented regardless of their color. Thus, counting the masks results in a significant error. Interestingly, as we can see, the false positive masks do not negatively affect LVLM-Count’s final answer, as the counting is done by an LVLM at the final stage, which is much stronger than GroundingDINO at understanding referring expressions. The only effect is that the white eggs have not been cut through by the division lines either. In the bottom row, we have chosen an image from the challenging Emoji-Count benchmark. The image contains icons, all of which have an arrow but point in different directions. However, the objects of interest are only “right arrows curving left.” Similar to the eggs example, taking the masks used for object-aware division results in a significant error.
@@ Appendix H Performance Analysis of LVLM-Count for different ground truth ranges on FSC-147 dataset
To further investigate the performance of our pipeline, we divide the ground truth values in the FSC-147 test set into intervals and plot the MAE for the base GPT-4o, Qwen2, and Gemma 3 models alongside the results from LVLM-Count using each model, as shown in Figure 20. The first interval contains relatively small ground truth values, a range in which LVLMs already perform well. As the ground truth values increase, the base models exhibit increasingly larger errors compared to LVLM-Count, with the margin growing rapidly. This behavior is consistent with our observations of counting errors on the blue circles in Figure 2.
@@
Figure 20: Performance analysis of our method, LVLM-Count, on the FSC-147 test set using GPT-4o (Figure 20(a)), Gemma 3 (Figure 20(b)), and Qwen2 (Figure 20(c)). In the first interval, all base LVLMs exhibit a lower MAE. However, in intervals with higher ground truth values, LVLM-Count achieves a lower MAE compared to the base LVLMs. Note that as the ground truth values grow higher, the margin of improvement increases rapidly.
@@
This section presents various accuracy measures for the experiments reported in Tables 1, 2, 3, and 4. The accuracy metrics are defined in Table 7. Following these definitions, Table 8 shows the measured accuracies for the FSC-147 test set. Similarly, Tables 9, 10, and 11 present the accuracy metrics for the PASCAL VOC, Emoji-Count, and Penguin benchmarks, respectively. The results in the tables demonstrate that LVLM-Count generally achieves a higher accuracy than the base LVLM it employs.
Metric Definition Acc Percentage of answers such that $answer=GT$ Acc±$\pm 1$ Percentage of answers such that $|answer-GT|\leq 1$ Acc±$\pm 3$ Percentage of answers such that $|answer-GT|\leq 3$ Acc±$\pm 5$ Percentage of answers such that $|answer-GT|\leq 5$ Acc±$\pm 10$ Percentage of answers such that $|answer-GT|\leq 10$
@@
Method Acc (%) Acc±$\pm$1 (%) Acc±$\pm$3 (%) Acc±$\pm$5 (%) Acc±$\pm$10 (%) GPT4o 12.24 26.22 42.10 52.41 66.58 LVLM-Count (GPT4o as LVLM) 14.45 ($\uparrow$) 28.99 ($\uparrow$) 47.25 ($\uparrow$) 57.03 ($\uparrow$) 70.45 ($\uparrow$) Gemma 3 27B 8.71 18.94 32.55 41.57 55.15 LVLM-Count (Gemma 3 as LVLM) 11.37 ($\uparrow$) 25.46 ($\uparrow$) 41.51 ($\uparrow$) 52.07 ($\uparrow$) 67.09 ($\uparrow$) Qwen2 VL 72B AWQ 9.19 20.81 36.83 46.95 62.07 LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 9.61 ($\uparrow$) 23.95 ($\uparrow$) 43.05 ($\uparrow$) 53.78 ($\uparrow$) 69.08 ($\uparrow$)
@@
Method Acc (%) Acc±$\pm$1 (%) Acc±$\pm$3 (%) Acc±$\pm$5 (%) Acc±$\pm$10 (%) GPT4o 30.39 47.06 61.11 71.90 89.87 LVLM-Count (GPT4o as LVLM) 32.35 ($\uparrow$) 51.96 ($\uparrow$) 73.20 ($\uparrow$) 82.35 ($\uparrow$) 93.46 ($\uparrow$) Gemma 3 27B 29.41 54.58 75.49 85.62 92.16 LVLM-Count (Gemma 3 27B as LVLM) 30.39 ($\uparrow$) 50.98 ($\downarrow$) 78.43 ($\uparrow$) 85.62 (-) 94.77 ($\uparrow$) Qwen2 VL 72B AWQ 24.18 42.48 60.13 72.22 86.60 LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 28.76 ($\uparrow$) 46.73 ($\uparrow$) 67.32 ($\uparrow$) 78.43 ($\uparrow$) 89.87 ($\uparrow$)
@@
Method Acc (%) Acc±$\pm$1 (%) Acc±$\pm$3 (%) Acc±$\pm$5 (%) Acc±$\pm$10 (%) GPT4o 1.85 5.54 13.73 21.12 43.94 LVLM-Count (GPT4o as LVLM) 2.81 ($\uparrow$) 9.88 ($\uparrow$) 23.53 ($\uparrow$) 36.22 ($\uparrow$) 60.24 ($\uparrow$) Gemma 3 27B 0.80 1.77 4.10 7.07 16.14 LVLM-Count (Gemma 3 27B as LVLM) 2.49 ($\uparrow$) 5.30 ($\uparrow$) 13.33 ($\uparrow$) 19.92 ($\uparrow$) 38.63 ($\uparrow$) Qwen2 VL 72B AWQ 0.88 2.89 7.55 11.41 19.76 LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 2.33 ($\uparrow$) 6.75 ($\uparrow$) 16.22 ($\uparrow$) 25.70 ($\uparrow$) 43.86 ($\uparrow$)
@@
Method Acc (%) Acc±$\pm$1 (%) Acc±$\pm$3 (%) Acc±$\pm$5 (%) Acc±$\pm$10 (%) GPT4o 2.00 3.67 8.33 12.00 22.00 LVLM-Count (GPT4o as LVLM) 2.67 ($\uparrow$) 6.33 ($\uparrow$) 14.33 ($\uparrow$) 18.67 ($\uparrow$) 33.00 ($\uparrow$) Gemma 3 27B 1.33 1.67 7.00 8.67 16.00 LVLM-Count (Gemma 3 27B as LVLM) 1.00 ($\downarrow$) 3.67 ($\uparrow$) 9.33 ($\uparrow$) 12.67 ($\uparrow$) 23.67 ($\uparrow$) Qwen2 VL 72B AWQ 1.67 2.00 6.33 10.33 16.67 LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 1.67 (-) 5.33 ($\uparrow$) 11.67 ($\uparrow$) 18.00 ($\uparrow$) 31.00 ($\uparrow$)

### v4

@@ 4.2 Results
Method MAE $\downarrow$ $\Delta$ MAE 95% CI RMSE $\downarrow$ $\Delta$ TFOC (Shi et al., 2024) 12.03 - - 18.18 - $\text{DAVE}_{\text{prm}}$ (Pelhan et al., 2024) 12.39 - - 22.81 - CountGD (Amini-Naieni et al., 2024) 2.81 - - 7.01 - GroundingRec (Dai et al., 2024) 4.05 - - 7.80 - Number of the target segmentation masks 4.03 - - 7.47 - GPT-4o counting target segmentaion masks 6.60 - - 15.09 - GPT4o 4.64 - [4.45, 4.82] 8.56 - LVLM-Count (GPT4o as LVLM) 3.42 $\downarrow$ 1.22 [3.25, 3.59] 7.17 $\downarrow$ 1.39 Gemma 3 27B 3.40 - [3.24 ,3.55] 7.54 - LVLM-Count (Gemma 3 27B as LVLM) 2.97 $\downarrow$ 0.43 [2.82, 3.11] 6.16 $\downarrow$ 1.38 Qwen2 VL 72B AWQ 4.80 - [4.61, 4.99] 8.71 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 4.12 $\downarrow$ 0.68 [3.87, 4.37] 7.76 $\downarrow$ 0.95
Emoji-Count. We evaluate the performance of the LVLMs and their performance using LVLM-Count on the Emoji-Count benchmark. The results are shown in Table˜3. This is a challenging benchmark, as it requires understanding complex concepts. We observe that taking the number of masks as the final count performs particularly poorly, as for any object of interest in the image, the segmentation stage tends to segment all the objects and cannot distinguish between different icons. Although GPT-4o and Gemma 3 show reasonable performance, the other open-source model does not perform well. Nonetheless, the performance of all three base LVLMs is significantly enhanced by LVLM-Count, especially Qwen2, which performs almost on par with GPT-4o when LVLM-Count is used for it.
@@
Method MAE $\downarrow$ $\Delta$ MAE 95% CI RMSE $\downarrow$ $\Delta$ TFOC (Shi et al., 2024) 64.64 - - 87.45 - $\text{DAVE}_{\text{prm}}$ (Pelhan et al., 2024) 198.99 - - 208.08 - CountGD (Amini-Naieni et al., 2024) 137.93 - - 156.80 - GroundingREC (Dai et al., 2024) 143.22 - - 158.74 - Number of the target segmentation masks 82.47 - - 107.98 - GPT-4o counting target segmentaion masks 107.72 - - 162.12 - GPT-4o 23.57 - [22.37, 24.76] 36.97 - LVLM-Count (GPT-4o as LVLM) 16.57 $\downarrow 7$ [16.25, 16.90] 33.11 $\downarrow 3.86$ Gemma 3 27B 21.39 - [21.32, 21.45] 24.04 - LVLM-Count (Gemma3 27B as LVLM) 16.16 $\downarrow 5.23$ [15.76, 16.56] 21.27 $\downarrow 2.77$ Qwen2 VL 72B AWQ 78.05 - [72.17, 83.93] 159.22 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 24.43 $\downarrow 53.62$ [23.65, 25.21] 43.38 $\downarrow 115.84$
Penguin Benchmark. We report the counting performance on the Penguin benchmark in Table˜4. We evaluate two variants of LVLM-Count. The primary variant uses GroundingDINO in both the area detection and target segmentation stages. The alternative variant does not use a detection model; instead, SAM is configured to segment all entities in the image. Both variants improve MAE across all three LVLMs, highlighting LVLM-Count’s robustness in scenarios with heavy occlusion and complex backgrounds—conditions that pose significant challenges for area detection and target segmentation. Additional details and a visual example showing the segmentation masks and division paths generated in both variants for a sample from the Penguin benchmark can be found in Appendix˜D.
@@
Method MAE $\downarrow$ $\Delta$ MAE 95% CI RMSE $\downarrow$ $\Delta$ GPT4o 35.18 - [33.13, 37.23] 45.76 - LVLM-Count( Main variant, using GPT-4o ) 26.76 $\downarrow$ 8.42 [25.86, 27.66] 38.60 $\downarrow$ 7.16 LVLM-Count(SAM-only variant, using GPT-4o) 29.02 $\downarrow$ 6.16 [28.21, 29.84] 44.89 $\downarrow$ 0.87 Gemma 3 27B 49.44 - [48.94, 49.95] 60.20 - LVLM-Count( Main variant, using Gemma 3 ) 34.13 $\downarrow$ 15.31 [31.72, 36.53] 44.11 $\downarrow$ 16.09 LVLM-Count( SAM-only variant, using Gemma 3 ) 41.09 $\downarrow$ 8.35 [39.92, 42.25] 53.01 $\downarrow$ 7.19 Qwen2 VL 72B AWQ 44.02 - [40.45, 47.60] 65.59 - LVLM-Count( Main variant, using Qwen2 ) 28.21 $\downarrow$ 15.81 [26.48, 29.94] 40.40 $\downarrow$ 25.19 LVLM-Count(SAM-only, using Qwen2 ) 33.33 $\downarrow$ 10.69 [30.45, 36.21] 47.55 $\downarrow$ 18.04
5 Limitations
@@
Acknowledgments
K. Fountoulakis would like to acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC). Cette recherche a été financée par le Conseil de recherches en sciences naturelles et en génie du Canada (CRSNG), [RGPIN-2019-04067, DGECR-2019-00147].
@@ Appendix A Ablation Study
We perform ablation studies using GPT-4o, conducting experiments on the FSC-147 test set. Table˜5 presents the results of these ablation experiments.
@@
Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ GPT-4o 25.57 - 137.26 - GPT-4o + Naive division 33.04 $\uparrow$ 7.47 116.83 $\downarrow$ 20.43 GPT-4o + Area Detection + Naive division 32.69 $\uparrow$ 7.12 102.41 $\downarrow$ 34.85 GPT-4o + Area detection 23.08 $\downarrow$ 2.49 120.06 $\downarrow$ 17.2 GPT-4o + Object-aware division (using SAM without GroundingDINO) 21.01 $\downarrow$ 4.56 135.04 $\downarrow$ 2.22 GPT-4o + Object-aware division 19.17 $\downarrow$ 6.40 120.61 $\downarrow$ 16.65 GPT-4o + Area detection + Object-aware division, (equiv. to LVLM-Count) 17.86 $\downarrow$ 7.71 91.71 $\downarrow$ 45.55
@@
Figure˜8 illustrates three examples of LVLM-Count’s performance on FSC-147. Additionally, Figure˜9 includes an example from the PASCAL VOC benchmark. Finally, we present three visual examples from the Emoji-Count benchmark in Figure˜10. In these visual examples, LVLM-Count achieves more accurate results compared to the base GPT-4o.
@@ Appendix C Robustness to Crowded Scenes and Heavy Occlusion
Figure˜11 demonstrates the strong performance of our method on a randomly selected image from the web containing various types of flowers. This scene features extreme occlusion; nonetheless, our method effectively handles the division task. It is important to note that the strong performance of our method in this challenging scenario is not coincidental. Figure˜12 illustrates the impact of our post-processing operations. The upper row displays the unprocessed SAM masks, while the bottom row shows the results after post-processing, which generates reliable paths for the division lines.
@@
To further demonstrate the effectiveness of our method in crowded scenes, we present Table˜6, which compares the performance of different LVLMs on FSC-147 dataset samples where the ground truth count is 100 or higher. The table contrasts their baseline performance with their performance when using our LVLM-Count method. The results show significant improvements when LVLM-Count is employed. Furthermore, Figure˜13 presents three visual examples of LVLM-Count’s performance on crowded images from FSC-147.
@@ Appendix D Additional Information and Visual Examples from the Penguin Benchmark
We construct our benchmark using the separate-site split. Given the dataset’s substantial size, which comprises tens of thousands of images, we randomly select 100 samples from the chosen split. To ensure balance, the sampling probability for an image with a specific ground-truth annotation count is inversely proportional to the frequency of that count in the entire split. We evaluate two variants of our pipeline. The main variant employs GroundingDINO with a low threshold for area detection and target segmentation, while the alternative variant does not use GroundingDINO at all. In this variant, masks are obtained by configuring SAM to segment any entity in the image. The results presented in Table˜4 show that both variants significantly improve the counting performance of the LVLMs, demonstrating the robustness of the initial stages in our pipeline.
Figure˜14 demonstrates the segmentation masks and division paths for the two variants on a sample from the Penguin benchmark. In Figure˜14 (a), GroundingDINO with a low detection threshold is employed. It correctly identifies all penguin instances, though there are some false positives. These detections are subsequently segmented by SAM. Figure˜14 (c) depicts the division paths found using these masks. In Figure˜14 (b), GroundingDINO is not used, and SAM is configured to segment all entities in the scene. As a result, in addition to the penguins, a large number of other objects are segmented. Nevertheless, Figure˜14 (d) shows that the division paths derived from these masks successfully partition the image without cutting any penguins. Note that some excessively large masks generated by SAM have been removed due to post-processing in our pipeline.
@@ Appendix E Real-world Application of LVLM-Count
As stated in Section˜1, counting has numerous real-world applications, including but not limited to biology, health, industry, warehousing, and environmental monitoring. Below, we demonstrate the performance of LVLM-Count on examples from the following areas: i) biology/health, ii) industry/warehousing, and iii) environmental monitoring. We also compare its results with those of the base LVLM (GPT-4o for the figures in this section). Note that in all examples, the cluster-based approach automatically determines the start and end points of the division paths.
In Figure˜15, images of two laboratory samples are analyzed using LVLM-Count. The first row shows an image from a dataset introduced by Lempitsky and Zisserman [2010], which contains simulated bacterial cells from fluorescence-light microscopy, created by Lehmussola et al. [2007]. The second row shows an image from the BM dataset introduced by Kainz et al. [2015], which contains bone marrow samples from eight healthy individuals. The standard staining procedure highlights the nuclei of various cell types in blue, while other cellular components appear in shades of pink and red [Paul Cohen et al., 2017]. As observed, LVLM-Count achieves much higher accuracy in counting bacterial cells and bone marrow nuclei in the top and bottom rows of Figure˜15, respectively, compared to the base LVLM, particularly for the bone marrow nuclei.
In Figure˜16, two images from industrial scenes are analyzed using LVLM-Count. The top row shows a sectional image of a stockpile of tree logs, and the bottom row shows an image from an industrial area containing barrels of various colors. For the top image, the objects of interest are the tree logs, while for the bottom image, LVLM-Count is tasked with counting the blue barrels. In both cases, LVLM-Count’s predictions are significantly closer to the ground truth values than those of the base LVLM, particularly for the tree logs, where the ground truth number is too large for the base LVLM to estimate accurately.
Figure˜17 shows an image sourced from a dataset [Penguin Research, 2016] created as part of an ongoing initiative to monitor the penguin population in Antarctica. This dataset comprises images captured hourly by a network of fixed cameras installed at over 40 locations. Over several years, this effort has accumulated over 500,000 images. Zoologists use these images to identify trends in penguin population sizes at each site, facilitating studies on potential correlations with factors such as climate change. Thus, determining the number of penguins in each image is crucial. Given the challenges of engaging human annotators to process such a vast dataset, automating the counting task is highly desirable [Arteta et al., 2016]. LVLM-Count is prompted to count the number of penguins in the image, and as observed, its predictions are significantly closer to the ground truth than those of the base LVLM.
@@ Appendix F LVLM-Count’s Power in Handling Multiple Object Categories in the Same Image
The image in Figure˜18 contains three object categories: person, cow, and horse. In the top row, the object of interest is “cow.” We prompt LVLM-Count to count the cows. First, the masks are produced through the initial stages of our pipeline, and then the cluster-based approach is used to automatically determine the start and end points of the division paths. It can be observed that horses have also been masked as cows. Nonetheless, this does not negatively impact the final answer; it merely causes the division lines to avoid cutting through the horses as well. The counting in LVLM-Count is performed by an LVLM (GPT-4o in this figure) and does not rely on the masks. We observe that GPT-4o successfully counts the number of cows in the resulting subimages, leading to the correct final answer.
In the middle row of Figure˜18, the object of interest is “person.” LVLM-Count again successfully counts the number of people accurately. A more interesting case is the bottom row of Figure˜18, where both cows and persons are objects of interest. We prompt LVLM-Count to count the number of “cows and persons.” Similar to the first row of the figure, there are false positive masks here as well. However, LVLM-Count successfully counts the number of instances from both categories combined since the counting is ultimately performed by the LVLM. Note that the number of objects in this image is limited, and GPT-4o might answer these questions correctly without the need for the LVLM-Count pipeline. This image has been chosen to illustrate LVLM-Count’s power in handling multiple objects in a counting task rather than for comparison with the baseline LVLM.
@@ Appendix G False Positive Masks at the Target Segmentation Stage
For example, in the top row of Figure˜19, E=E=brown egg”. However, all the eggs have been segmented regardless of their color. Thus, counting the masks results in a significant error. Interestingly, as we can see, the false positive masks do not negatively affect LVLM-Count’s final answer, as the counting is done by an LVLM at the final stage, which is much stronger than GroundingDINO at understanding referring expressions. The only effect is that the white eggs have not been cut through by the division lines either. In the bottom row, we have chosen an image from the challenging Emoji-Count benchmark. The image contains icons, all of which have an arrow but point in different directions. However, the objects of interest are only “right arrows curving left.” Similar to the eggs example, taking the masks used for object-aware division results in a significant error.
@@ Appendix H Performance Analysis of LVLM-Count for different ground truth ranges on FSC-147 dataset
To further investigate the performance of our pipeline, we divide the ground truth values in the FSC-147 test set into intervals and plot the MAE for the base GPT-4o, Qwen2, and Gemma 3 models alongside the results from LVLM-Count using each model, as shown in Figure˜20. The first interval contains relatively small ground truth values, a range in which LVLMs already perform well. As the ground truth values increase, the base models exhibit increasingly larger errors compared to LVLM-Count, with the margin growing rapidly. This behavior is consistent with our observations of counting errors on the blue circles in Figure˜2.
@@
Figure 20: Performance analysis of our method, LVLM-Count, on the FSC-147 test set using GPT-4o (Figure˜20(a)), Gemma 3 (Figure˜20(b)), and Qwen2 (Figure˜20(c)). In the first interval, all base LVLMs exhibit a lower MAE. However, in intervals with higher ground truth values, LVLM-Count achieves a lower MAE compared to the base LVLMs. Note that as the ground truth values grow higher, the margin of improvement increases rapidly.
@@
This section presents various accuracy measures for the experiments reported in Tables 1, 2, 3, and 4. The accuracy metrics are defined in Table˜7. Following these definitions, Table˜8 shows the measured accuracies for the FSC-147 test set. Similarly, Tables 9, 10, and 11 present the accuracy metrics for the PASCAL VOC, Emoji-Count, and Penguin benchmarks, respectively. The results in the tables demonstrate that LVLM-Count generally achieves a higher accuracy than the base LVLM it employs.
@@
[…]
@@
[…]
@@
[…]
@@
[…]